A sphere of mass 20 kg is attracted by another sphere of mass

30 kg with a force of 4x 10-5 V. What is the distance between the

spheres? Round your answer to the nearest hundredth.


1
Expert's answer
2022-04-27T13:33:57-0400
"F=G\frac{m_1m_2}{r^2}"

"r=\sqrt{G\frac{m_1m_2}{F}}"

"r=\sqrt{6.67*10^{-11}\frac{20*30}{4*10^{-5}}}=0.032\:\rm m=3.2\: cm"